OpenStudy (anonymous):

find the coordinates of the midpoint of a segment whose endpoints are (-4 6) and (-8 -2)

OpenStudy (anonymous):

\[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\] midpoint formula basically taking the average of both the coordinates.